OpenStudy (anonymous):

In a dog show, 20% of the dogs are eliminated each day. If the competition starts with 2,000 dogs, to the nearest integer, how many dogs will be present at the start of the 5th day? a. 259 b. 480 c. 819 d. 1,229 Confused....

OpenStudy (anonymous):

by eliminated do you mean killed? because that makes this question so much more exciting!

OpenStudy (anonymous):

haha...

OpenStudy (anonymous):

so we have a geometric sequence, with the initial value = 2000 common ratio = 100% - 20% = 80% = 0.8 Un = ar^(n-1) n=5 in this case \[2000 \times (0.8)^4\]

OpenStudy (anonymous):

have you done geometric sequences before?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so wait u have to subtract 100-20?

OpenStudy (anonymous):

so if we do that tahn its 819.2 so c?

OpenStudy (anonymous):

\[2000 \times (1 \div 5)^{4}\] since its the 5th day and the dogs haven't been eliminated on the fifth day

OpenStudy (anonymous):

oh damn \[2000 \times (4 \div 5)^4\] i din't read the question carefully lol

OpenStudy (anonymous):

is ti c?